OpenStudy (anonymous):

How do I kno how to solve this equation either by substitution or elimination 3x = -31 +2y and 5x + 6y = 23? Please help me and explain how

OpenStudy (anonymous):

Type in Wolfram Alpha into Google and put in the problem into the box

OpenStudy (anonymous):

Are u serious or kidding me?

OpenStudy (texaschic101):

I would solve by elimination. rearrange the first equation... 3x = -31 + 2y is the same as 3x - 2y = -31 5x + 6y = 23 so now we have : 3x - 2y = -31 -->(3) 5x + 6y = 23 --------------- 9x - 6y = - 93 (result of multiplying by 3) 5x + 6y = 23 ---------------add 14x = - 70 x = - 5 now sub -5 in for x in either of the original equations 3x - 2y = -31 3(-5) - 2y = -31 -15 - 2y = -31 -2y = -31 + 15 -2y = - 16 y = 8 check... 5x + 6y = 23 5(-5) + 6(8) = 23 -25 + 48 = 23 23 = 23 (correct) x = -5 and y = 8 any questions ?

OpenStudy (anonymous):

No thanx!!! :)

OpenStudy (texaschic101):

glad I could help :)